OpenStudy (anonymous):

Let y=3x^2 Find the change in y, ∆y when x=4 and ∆x=0.2_____? Find the differential dy when x=4 and dx=0.2______?

jimthompson5910 (jim_thompson5910):

When x = 4, y = 3(4)^2 = 3(16) = 48 Now add on 0.2, to get x = 4.2 and y = 3(4.2)^2 = 3(17.64) = 52.92 Notice that y goes from 48 to 52.92. So the change in y is 52.92 - 48 = 4.92 This means that ∆y = 4.92

OpenStudy (anonymous):

How can I find the other one?

jimthompson5910 (jim_thompson5910):

use the idea that dy/dx = derivative of y with respect to x So basically derive y = 3x^2 and you'll get something in the form dy/dx = f(x), where f(x) is your derivative function From there, plug in dx = 0.2 and solve for dy

jimthompson5910 (jim_thompson5910):

Oh plug in x = 4 as well into the function when you plug in dx = 0.2